 |
 |
- 2001323+1411 (note, jrizzo)
-
This DCP has been removed and will be relocated.
- 2001323+1410 (retired, jrizzo)
-
Data for this station is downloaded every 24 hours.
- 2001323+1410 (retired, jrizzo)
-
head vln tilt roll vlu vle :
No data from the current meter again. Can't communicate with the RDI directly through cube.
- 2001279+1424 (retired, jrizzo)
-
vlu vle :
No data from the current meter since day 274. I've logged in to the system and talked to the cube. No RDI data was being recorded. I talked to the RDI just fine manually, then I watched the cube take 2 automated measurements. We'll have to monitor the system
- 2001268+2238 (retired, duff)
-
No data retrieved since 2001263. The phone line and/or modem is somehow hosed again. Will investigate. Modem communication re-established. Data downloaded.
- 2001264+2047 (retired, duff)
-
all :
2001248 (09/05/2001 00:00 - 09/06/2001 00:00) :
There is a gap for all of the sensors on day 248. This data will not be recovered. No active message required.
- 2001264+1429 (retired, jrizzo)
-
No data recieved since day 259. Data are being retrieved once again after crew reset PC at Freeport office.
- 2001249+1429 (retired, duff)
-
No data from current meter since day 244. John Adams made a site visit, switched the serial port that the current meter uses and we are now collecting data.
- 2001236+1456 (retired, duff)
-
No data downloaded since 2001232+2300. Fixed.
- 2001197+1756 (retired, jrizzo)
-
1/16/2001 (01/16/2001 00:00 - 01/17/2001 00:00) :
Rainy weather preventing reinstallation of cube and batteries. Cube reinstalled
- 2001197+1755 (retired, jrizzo)
-
all :
2001026+0930 (01/26/2001 09:30 - 01/26/2001 09:31) :
RDI current meter has been removed to CBI lab for troubleshooting. RDI has been reinstalled
- 2001081+1232 (retired, jrizzo)
-
today (05/12/2025 00:00 - 05/13/2025 00:00) :
System reinstalled, but no communication with the RDI. System removed
- 2001054+1427 (retired, terry)
-
No current data since 348+0554. RDI has been removed for repair 2001026.
- 2000347+1616 (retired, jrizzo)
-
No RDI data being recieved. RDI has been reactivated.
- 2000344+1335 (retired, jrizzo)
-
No current data received since 2000340+2118. Batteries replaced and cube reset. Scott verified operation before field crew left the site.
- 2000336+1540 (retired, ussery)
-
vln vlu vle :
No current data received since 332+1448. Data received since 334+1612.
- 2000325+2025 (retired, jrizzo)
-
no data since 2000323. Station dies when battery reached 7.8v. Battery "DEAD" Batteries replaced and system is operational.
- 2000319+1329 (retired, duff)
-
No data received since 314+0606. Power problems with the computer at the Port Office. Went through a debugging session with a secretary named Jan, to no avail. Computer has no power, but UPS is on. ***Field crew made a site visit and replaced UPS. System is up ***
- 2000311+1809 (retired, duff)
-
No data downloaded since 304+0600.
The /var partition on the fredow cube filled up, thus causing the dlmark file to become empty, thus causing the data collection computer to download everything from the cube each time it connects. This means that the daily logs on the data collection computer are in the tens of megabytes. The problem with the cube has been fixed, but I don't know what to do about filling in the data locally yet. (I'm not about to transfer approximately 119 megs for three days worth of data over a 14.4 modem connection)
***Data recovered***
- 2000283+1454 (retired, jrizzo)
-
con :
No depth being reported
|
 |
 |